Suppose you lived in a pre-industrial society and needed to lift a heavy (20 kg) block a height of 5 m and had two choices for how to accomplish your task:
1
Expert's answer
2020-05-04T12:46:24-0400

Let's break the question into two parts:


1) The force needed in Ramp scenario.

2) The effort force needed in the lever scenario.


1. Ramp Scenario: 

In an incline, the only component of cart's weight that is in the direction of motion is "mg\sin{\alpha}" . Therefore the effort force in this case must be equal or greater than "mg\sin{\alpha}".

Now we need to find the angle between the incline of the ramp and the ground. 

Since the height is 5m and the length of the ramp is 8m, would be "\frac{5}{8}" or 0.625.


"F=0.625(20)(10)=125\ N"

Therefore, the minimum Effort force you would require in this case is 125N.


2. Lever Scenario:

Just apply "moment action" in this case, which is:

 

"F_ed_e=F_rd_r"

"F_r=20(10)=200\ N, d_e=10\ m, F_r=1\ m"

Plug-in the values in the above equation:

 

"F_e=\frac{200}{10}= 20N"

As 20 N < 125 N, the best choice is to use lever.